# Neural Conditional Transport Maps

### v1

@@ Abstract
We present a neural framework for learning conditional optimal transport (OT) maps between probability distributions. Our approach introduces a conditioning mechanism capable of processing both categorical and continuous conditioning variables simultaneously. At the core of our method lies a hypernetwork that generates transport layer parameters based on these inputs, creating adaptive mappings that outperform simpler conditioning methods. Comprehensive ablation studies demonstrate the superior performance of our method over baseline configurations. Furthermore, we showcase an application to global sensitivity analysis, offering high performance in computing OT-based sensitivity indices. This work advances the state-of-the-art in conditional optimal transport, enabling broader application of optimal transport principles to complex, high-dimensional domains such as generative modeling and black-box model explainability.
@@ 1 Introduction
Optimal transport (OT) is a powerful mathematical framework for comparing and transforming probability distributions, with wide applications across machine learning, computer vision, and scientific computing. OT provides a natural geometry for distribution spaces, offering stronger theoretical guarantees compared to alternative measures (Peyré and Cuturi, 2019). Despite its theoretical appeal, applying OT to high-dimensional, real-world problems has long been constrained by computational limitations. Classical OT methods scale poorly with dimensionality and sample size. Neural approaches have made significant progress in addressing these challenges by approximating transport maps with neural networks (Korotin et al., 2023; Makkuva et al., 2020), enabling efficient computation in high-dimensional spaces.
@@
However, many practical applications require conditional OT maps—transformations that adapt based on auxiliary variables such as labels, time indices, or other parameters. For instance, in climate-economy models, we need to model how distributions of climate variables change based on emissions or policy scenarios. This capability is essential for emulators of computationally intensive models for comprehensive uncertainty quantification. The challenge lies in efficiently computing these conditional transport maps, particularly in data-intensive problems where traditional OT methods become computationally prohibitive.
@@
Global sensitivity analysis (GSA) is another compelling application for conditional OT. GSA quantifies how uncertainty in model outputs can be attributed to different input sources—critical for understanding complex black-box models in climate science, economics, and machine learning (Borgonovo et al., 2024). Recent works leverage OT costs to define GSA indices (Wiesel, 2022; Borgonovo et al., 2024), offering valuable theoretical properties. However, these methods remain constrained by the computational scalability of the underlying OT solvers, limiting their applicability to real-world scientific questions. Efficiently learning conditional transport maps can therefore enable both more robust uncertainty-aware generative models and provide new tools for large-scale black-box model explainability.
@@
In this paper, we introduce a neural framework that efficiently learns conditional OT maps across both categorical and continuous conditioning variables. Our approach leverages a hypernetwork architecture—a neural network that dynamically generates the parameters for transport layers based on conditioning inputs. This mechanism creates highly adaptive mappings that significantly outperform simpler conditioning methods. Our contributions are as follows:
@@
  • We introduce a conditioning mechanism capable of simultaneously processing both categorical and continuous variables, using learnable embeddings and positional encoding.
  • We propose a hypernetwork-based architecture that generates condition-specific transformation parameters, enabling fundamentally different mappings for each condition value.
  • We provide extensive empirical validation across synthetic datasets, climate data, and integrated assessment models, demonstrating superior performance compared to baselines.
  • We show an application to global sensitivity analysis, enabling efficient black-box model.
  • Upon publication, we will release an open-source implementation of our method and data.
The remainder of this paper is organized as follows. Section 2 reviews related work on traditional and neural OT, and conditioning methods. Section 3 details our conditional neural transport framework, including problem formulation, architecture, and training procedure. Section 4 presents results on benchmark datasets and ablations. Finally, Section 5 discusses limitations and future directions.
@@ 2 Background
Optimal Transport has emerged as a powerful mathematical framework across numerous domains. In machine learning, OT has been used for generative modeling (Arjovsky et al., 2017), domain adaptation (Courty et al., 2017), and representation learning (Tolstikhin et al., 2018). In graphics, OT has become fundamental for geometry processing (Solomon et al., 2015), point clouds (Bonneel et al., 2016), noise generation (De Goes et al., 2012), and appearance transfer (Pitié et al., 2007). In computer vision, OT enables image color transfer (Ferradans et al., 2014), shape analysis (Solomon et al., 2016), and texture synthesis (Gao et al., 2019). OT has also been applied to GSA (Borgonovo et al., 2024; Wiesel, 2022), offering sensitivity indices with strong statistical properties.
Neural Optimal Transport approaches began with Wasserstein GANs (Arjovsky et al., 2017), which approximate Wasserstein distances without computing explicit transport maps. Later methods based on Brenier’s theorem (Brenier, 1991) used input convex neural networks (ICNNs) (Makkuva et al., 2020; Amos et al., 2017) to implement Monge maps. However, this formulation is limited to a specific ground cost ($L_{2}^{2})$, and by the existence of Monge maps. Moreover, ICNNs impose severe architectural constraints—requiring non-negative weights, limited activation functions, and specialized initialization—leading to reduced expressivity and optimization difficulties, particularly for conditional problems Korotin et al. (2021). More recently, Korotin et al. introduced neural optimal transport (NOT) (Korotin et al., 2023) and Kernel NOT (Korotin et al., 2022), bypassing these limitations through a minimax formulation supporting general cost functions and stochastic maps. For conditional OT, existing approaches either rely on these restrictive ICNNs (CondOT (Bunne et al., 2022)) or other architectural constraints (Wang et al., 2023). Notably, these methods do not provide standardized benchmarking datasets, hindering systematic comparison across different conditional transport approaches. Our work extends NOT to conditional settings with flexible conditioning mechanisms for both discrete and continuous variables, avoiding the expressivity constraints and training difficulties associated with previous approaches.
Conditioning mechanisms in generative models have evolved from simple concatenation (Ho et al., 2020; Mirza and Osindero, 2014) to more sophisticated approaches including classifier guidance (Dhariwal and Nichol, 2021), normalization (Huang and Belongie, 2017; Perez et al., 2018), attention (Vaswani et al., 2017; Rombach et al., 2022), and hypernetworks (Ha et al., 2017). These methods create a spectrum of expressivity, with hypernetworks offering greater flexibility by dynamically generating parameters for condition-specific transformations. For conditional transport maps, existing approaches like CondOT (Bunne et al., 2022) rely on simple concatenation, limiting their ability to model divergent transport behaviors. Our hypernetwork-based approach enables distinct transformations for different conditions, crucial when conditional distributions require fundamentally different mapping strategies. We provide the first quantitative comparison of conditioning mechanisms in neural optimal transport, showcasing the expressiveness of hypernetworks in this setting.
@@ 3 Neural Conditional Transport Maps
Conditional OT seeks to learn OT maps that can adapt to different conditions or contexts, an important capability for applications ranging from sensitivity analysis to conditional generative modeling. While classical OT methods struggle with computational scalability or conditioning flexibility, neural approaches offer a promising alternative. In this section, we present our framework for conditional neural OT, which extends the NOT approach of Korotin et al. (2023) with a hypernetwork-based conditioning mechanism. We illustrate our OT maps in Figure 1. We begin by formalizing the conditional OT problem (Section 3.1), then describe our neural architecture that leverages encoder-decoder structures for both transport map $T$ and critic function $f$ (Section 3.2). We introduce our conditioning mechanism that handles both discrete variables and continuous partitions (Section 3.3), present architectural variants including hypernetwork and self-attention approaches (Section 3.4), and detail our training procedure with a custom pretraining strategy (Section 3.5). Implementation details are provided in the supplementary material.
@@ 3.1 Problem formulation
Let us consider two probability measures $\mu$ and $\nu$ defined over two Polish spaces, $\mathcal{X}\subset\mathbb{R}^{m}$ and $\mathcal{Y}\subset\mathbb{R}^{n}$, respectively. Let’s also define a lower-semicontinuous cost function $k\colon\mathcal{X}\times\mathcal{Y}\longrightarrow[0,+\infty]$ such that $k(y,y^{\prime})=0$ if and only if $y=y^{\prime}$. The Kantorovich formulation of the OT problem can be stated as follows:
[…]
@@
Even though the primal formulation of the OT problem in Equation (1) is easier to interpret, the dual formulation is more useful from a practical perspective, especially using neural networks. First, we introduce the concept of $k$-transform. The $k$-transform $f^{k}$ of a function $f\colon\mathcal{Y}\longrightarrow\mathbb{R}$ is defined as $f^{k}(x)\coloneq\inf_{y\in\mathcal{Y}}\{k(x,y)-f(y)\}$. The dual formulation of the Kantorovich OT problem is then:
[…]
For our purposes, the dual problem can be further reformulated as a maximin problem (Korotin et al., 2023). We introduce a third atomless distribution, $\omega$, defined over the Polish space $\mathcal{Z}\subset\mathbb{R}^{s}$. Given a measurable map $T\colon\mathcal{X}\times\mathcal{Z}\longrightarrow\mathcal{Y}$ and its push-forward operator $T\#$, the maximin formulation is:
$\footnotesize K(\mu,\nu)=\sup_{f}\inf_{T}\mathcal{L}(f,T),$ (3)
@@
[…]
We note here that the results in Korotin et al. (2023) can be extended to weak costs, but they are out of the scope of this work.
The conditioned problem. We start from Equation (4) to integrate the conditioning mechanism into the problem formulation. We assume that the condition $c$ belongs to a measure space $\mathcal{C}$. From the theoretical perspective, the extension is straightforward:
@@
[…]

### v2

@@ Abstract
We present a neural framework for learning conditional optimal transport (OT) maps between probability distributions. Conditional OT maps are transformations that adapt based on auxiliary variables, such as labels, time indices, or other parameters. They are essential for applications ranging from generative modeling to uncertainty quantification of black-box models. However, existing methods for generating conditional OT maps face significant limitations: input convex neural networks (ICNNs) impose severe architectural constraints that limit expressivity. At the same time, simpler conditioning strategies, such as concatenation, fail to model fundamentally different transport behaviors across conditions. Our approach introduces a conditioning mechanism capable of simultaneously processing both categorical and continuous conditioning variables, using learnable embeddings and positional encoding. At the core of our method lies a hypernetwork that generates transport layer parameters based on these inputs, creating adaptive mappings that outperform simpler conditioning methods. We showcase the framework’s practical impact through applications to global sensitivity analysis, enabling efficient computation of OT-based sensitivity indices for complex black-box models. This work advances the state-of-the-art in conditional optimal transport, enabling broader application of optimal transport principles to complex, high-dimensional domains such as generative modeling, black-box model explainability, and scientific computing.
@@ 1 Introduction
Optimal transport (OT) is a powerful mathematical framework for comparing and transforming probability distributions, with wide applications across machine learning, computer vision, and scientific computing. OT provides a natural geometry for distribution spaces, offering stronger theoretical guarantees compared to alternative measures (Peyré and Cuturi, 2019). The importance of OT in machine learning has grown significantly, particularly in generative modeling where it forms the foundation of flow matching (Lipman et al., 2023) and rectified flows (Liu et al., 2023). Despite its theoretical appeal and practical success, applying OT to high-dimensional, real-world problems has long been constrained by computational limitations. Classical OT methods scale poorly with sample size. Neural approaches have made significant progress in addressing these challenges by approximating transport maps with neural networks (Arjovsky et al., 2017; Makkuva et al., 2020; Korotin et al., 2023), enabling efficient computation in high-dimensional spaces.
@@
The optimal mapping between distributions and the related cost are two elements of interest in the OT framework. The optimal mapping is used in different fields, ranging from computer vision (Bonneel and Digne, 2023) to biology (Bunne et al., 2023). However, many practical applications require conditional optimal maps: transformations that adapt based on auxiliary variables such as labels, time indices, or other parameters. Conditional transformations in general are crucial across diverse domains: generative models need to produce samples conditioned on class labels or continuous attributes, domain adaptation requires transport maps that vary with domain-specific characteristics, and climate-economy models need to model how distributions of climate variables change based on emissions or policy scenarios. The challenge lies in efficiently computing these conditional optimal maps, particularly in data-intensive problems where traditional OT methods become computationally prohibitive.
@@
The OT cost has recently been presented as a valuable tool for global sensitivity analysis (GSA, Wiesel (2022); Borgonovo et al. (2024)). This application exemplifies the computational challenge for conditional OT. GSA leverages statistical methods to quantify how uncertainty in model outputs can be attributed to different input sources (Saltelli, 2002). GSA is therefore useful for understanding complex black-box models in machine learning, climate science, and economics (Saltelli et al., 2020). Recent works leverage OT costs to define sensitivity indices with valuable theoretical properties (Wiesel, 2022; Borgonovo et al., 2024). However, these methods face a severe computational bottleneck: computing sensitivity indices based on OT for complex models with many inputs requires solving hundreds to thousands of OT problems. The classical approach becomes computationally intractable due to memory requirements for storing cost matrices and the combinatorial explosion of required OT solvers. Efficiently learning conditional transport maps provides new tools for large-scale black-box model explainability.
Existing neural conditional OT methods suffer from fundamental limitations that prevent deployment in general cases (Bunne et al., 2022; Wang et al., 2025). The first approach is based on input convex neural networks (ICNNs), imposing severe architectural constraints, requiring non-negative weights, limited activation functions, and specialized initialization, leading to reduced expressivity and optimization difficulties, particularly for complex conditional problems (Korotin et al., 2021). Moreover, using concatenation as a conditioning strategy fails to model complex transport patterns across conditions. This type of conditioning lacks the flexibility to learn distinct mappings when different conditions require qualitatively different transformations, limiting its applicability to real-world scenarios where conditioning variables significantly alter the underlying transport structure. The second approach, based on conditional OT flows, alleviates some of these expressivity issues by conditioning the velocity field of an ODE on auxiliary variables, allowing the entire transport trajectory to depend on the condition. However, this comes at the cost of more complex training: solving ODE constraints requires additional regularization terms, hyperparameter tuning is delicate, and optimization is significantly more expensive, making it difficult to scale to large datasets or high-dimensional problems. Finally, Wang et al. (2025) restricts attention to simple tractable input distributions, while we consider the general setting where input distributions are arbitrary.
@@
In this paper, we expand the state-of-the-art neural OT framework (Korotin et al., 2023) to efficiently learn conditional OT maps across both categorical and continuous conditioning variables. Our approach leverages a hypernetwork architecture—a neural network that dynamically generates the parameters for transport layers based on conditioning inputs. This mechanism creates adaptive mappings outperforming simpler conditioning methods while maintaining computational efficiency. Our contributions are as follows:
@@
  • We introduce a unified conditioning mechanism capable of simultaneously processing both categorical and continuous variables, using learnable embeddings and positional encoding.
  • After testing a variety of possible alternatives, we propose a hypernetwork-based architecture that generates condition-specific transformation parameters, enabling fundamentally different mappings for each condition value without the computational overhead of separate networks.
  • We provide empirical validation across synthetic datasets, image generation tasks, climate data, and integrated assessment models.
  • We demonstrate practical applications to global sensitivity analysis, enabling efficient computation of OT-based sensitivity indices for complex black-box models at scale. This is the first comparison of the neural OT framework by Korotin et al. (2023) with well-established solvers for this task.
  • We release our data and open-source implementation of our method in this Github repository.
The remainder of this paper is organized as follows. Section 2 reviews related work on traditional and neural OT, and conditioning methods. Section 3 details our conditional neural transport framework, including problem formulation, architecture, and training procedure. Section 4 presents the applications and experimental setup, while Section 5 shows results on benchmark datasets and ablations. Finally, Section 6 discusses limitations and future directions.
@@ 2.1 Optimal Transport Foundations and Modern Applications
Optimal transport (OT) has emerged as a powerful mathematical framework across numerous domains, providing principled methods for comparing and transforming probability distributions. In machine learning, OT has been applied to generative modeling (Arjovsky et al., 2017), domain adaptation (Courty et al., 2017), and representation learning (Tolstikhin et al., 2018). Beyond traditional applications, OT has become foundational to modern generative modeling paradigms: flow matching (Lipman et al., 2023) uses optimal transport displacement interpolation to define probability paths for training continuous normalizing flows, while rectified flows (Liu et al., 2023) employ OT principles to learn straight transport paths for efficient generative modeling. In graphics and computer vision, OT enables geometry processing (Solomon et al., 2015), point cloud analysis (Bonneel et al., 2016), image color transfer (Ferradans et al., 2014), and texture synthesis (Gao et al., 2019).
The growing importance of OT in scientific computing is demonstrated by its application to GSA (Wiesel, 2022; Borgonovo et al., 2024), where OT-based sensitivity indices offer valuable theoretical properties for understanding complex black-box models. However, classical OT methods face severe computational limitations when applied to real-world problems, motivating the development of neural approaches that can scale to modern machine learning applications.
2.2 Neural Optimal Transport
Neural optimal transport methods were initially developed to address the computational bottlenecks of classical OT solvers. Early approaches include Wasserstein GANs (Arjovsky et al., 2017), which approximate Wasserstein distances through adversarial training but do not compute explicit transport maps. Later methods leveraged Brenier’s theorem (Brenier, 1991) to implement Monge maps using input convex neural networks (ICNNs) (Amos et al., 2017; Makkuva et al., 2020).
However, ICNN-based approaches suffer from fundamental limitations that restrict their applicability. The convexity requirement imposes strict architectural constraints: all weights must be non-negative, activation functions are limited to convex choices (typically ReLU), and specialized initialization procedures are required to maintain convexity during training. More recently, Korotin et al. (2023) introduced Neural Optimal Transport (NOT) and Kernel NOT (Korotin et al., 2022), which bypass ICNN limitations through a minimax formulation that supports general cost functions and stochastic transport maps. NOT uses unconstrained neural networks for transport maps and critic functions, enabling flexible architectures while maintaining theoretical guarantees. We build upon this work for our conditional OT formulation.
2.3 Conditional Transport Challenges
Computational Bottlenecks: Traditional approaches to conditional transport problems require solving multiple independent OT problems for different conditioning values. This becomes computationally prohibitive at scale—for instance, global sensitivity analysis applications require computing OT-based sensitivity indices by solving hundreds to thousands of individual OT problems for each input variable and partition (Borgonovo et al., 2024). This approach is intractable for high-dimensional models with large datasets due to memory requirements for storing cost matrices and the combinatorial explosion of required solvers.
Methodological Limitations: Existing conditional OT methods suffer from architectural constraints that prevent effective scaling. CondOT (Bunne et al., 2022) relies on ICNNs, inheriting all the expressivity and optimization problems discussed above, which are further amplified in conditional settings where different conditions may require qualitatively different transport behaviors. Alternative approaches (Wang et al., 2023) impose architectural restrictions and assumptions on the input distributions that limit their flexibility. Furthermore, these methods typically use simple conditioning strategies like concatenation, which cannot model the different transport behaviors required when conditioning variables significantly alter the underlying distribution structure.
Benchmarking Limitations: Unlike other areas of machine learning, conditional optimal transport lacks standardized benchmarking and evaluation protocols. This absence hinders systematic comparison across different approaches and limits progress in the field. Existing methods often evaluate on different datasets with varying experimental setups, making it difficult to assess their relative strengths and weaknesses. Moreover, existing approaches focus on implementing the neural OT solver, while we also test various conditioning architectures.
2.4 Conditioning Mechanisms in Generative Models
The design of effective conditioning mechanisms is crucial for adaptive generative models. Conditioning approaches have evolved significantly, creating a spectrum of expressivity and computational complexity. Simple concatenation (Ho et al., 2020; Mirza and Osindero, 2014) directly combines feature vectors with conditioning information but lacks the flexibility to model condition-dependent behaviors. More sophisticated approaches include classifier guidance (Dhariwal and Nichol, 2021), which uses external classifiers to guide generation; normalization-based methods like FiLM (Perez et al., 2018) and AdaIN (Huang and Belongie, 2017), which modulate feature statistics; and attention mechanisms (Vaswani et al., 2017; Rombach et al., 2022), which allow selective focus on relevant conditioning information.
At the high end of the expressivity spectrum, hypernetworks (Ha et al., 2017) dynamically generate network parameters based on conditioning inputs, enabling fundamentally different computations for different conditions. This capability is particularly valuable for OT applications, where different conditioning values may require distinct mapping strategies that cannot be achieved through feature modulation alone.
For conditional transport maps specifically, the choice of conditioning mechanism is critical. Simple concatenation, as used in CondOT (Bunne et al., 2022), fails when different conditions require qualitatively different transport behaviors—for example, when transporting between fundamentally different distribution types or when conditioning variables induce non-linear changes in the optimal transport structure. Our hypernetwork-based approach addresses this limitation by generating condition-specific transformation parameters, providing the expressiveness of separate networks per condition while maintaining computational efficiency through parameter sharing in the hypernetwork itself.
This work provides the first systematic comparison of conditioning mechanisms within the neural optimal transport framework, demonstrating the superior performance of hypernetwork-based approaches for learning flexible conditional transport maps.
@@ 3 Neural Conditional Transport Maps
Our solution extends the Neural Optimal Transport (NOT) framework of Korotin et al. (2023) with a flexible conditioning architecture that enables efficient learning of conditional transport maps across both categorical and continuous conditioning variables. At the core of our method lies a hypernetwork that dynamically generates transport layer parameters based on conditioning inputs, providing high expressivity while maintaining computational efficiency through shared parameter generation. This design allows different conditions to induce different mapping strategies—essential when conditional distributions require distinct transport behaviors that cannot be captured through simple feature concatenation or modulation approaches.
Our framework offers several key advantages over existing methods: (1) it avoids the architectural constraints of ICNN approaches, enabling flexible network designs with standard optimization procedures; (2) it scales efficiently to large numbers of conditioning values without requiring separate OT solvers, addressing the computational bottlenecks in sensitivity analysis applications; (3) it provides a unified mechanism for handling both discrete and continuous conditioning variables through learnable embeddings and positional encoding; and (4) it enables systematic comparison of conditioning strategies within the same neural OT formulation.
We illustrate our conditional OT maps in Figure 1. In this section, we begin by formalizing the conditional OT problem (Section 3.1), then describe our architecture that leverages encoder-decoder structures for both transport map $T$ and critic function $f$ (Section 3.2). We then introduce our conditioning mechanism that handles discrete variables and continuous partitions (Section 3.3), present architectural variants including our hypernetwork approach and alternative conditioning strategies (Section 3.4), and detail our training procedure with a custom pretraining strategy that addresses the optimization challenges common in conditional transport learning (Section 3.5). Implementation details are provided in the supplementary material.
@@ 3.1 Problem formulation
To address the computational bottlenecks in conditional OT applications, we formulate the problem following the neural OT framework that enables efficient scaling to large numbers of conditioning values without requiring separate solvers for each condition. Let us consider two probability measures $\mu$ and $\nu$ defined over two Polish spaces, $\mathcal{X}\subset\mathbb{R}^{m}$ and $\mathcal{Y}\subset\mathbb{R}^{n}$, respectively. Let’s also define a lower-semicontinuous cost function $k\colon\mathcal{X}\times\mathcal{Y}\longrightarrow[0,+\infty]$ such that $k(y,y^{\prime})=0$ if and only if $y=y^{\prime}$. The Kantorovich formulation of the OT problem can be stated as follows:
$K(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}k(x,y)\,d\pi(x,y),$ (1)
@@
Even though the primal formulation of the OT problem in equation 1 is easier to interpret, the dual formulation is more useful from a practical perspective, especially using neural networks. First, we introduce the concept of $k$-transform. The $k$-transform $f^{k}$ of a function $f\colon\mathcal{Y}\longrightarrow\mathbb{R}$ is defined as $f^{k}(x)\coloneq\inf_{y\in\mathcal{Y}}\{k(x,y)-f(y)\}$. The dual formulation of the Kantorovich OT problem is then:
$K(\mu,\nu)=\sup_{f}\left[\int_{\mathcal{X}}f^{k}(x)d\mu(x)+\int_{\mathcal{Y}}f(y)d\nu(y)\right].$ (2)
For our purposes, the dual problem can be further reformulated as a maximin problem (Korotin et al., 2023). We introduce a third atomless distribution, $\omega$, defined over the Polish space $\mathcal{Z}\subset\mathbb{R}^{s}$. Given a measurable map $T\colon\mathcal{X}\times\mathcal{Z}\longrightarrow\mathcal{Y}$ and its push-forward operator T#T\#, the maximin formulation is:
$K(\mu,\nu)=\sup_{f}\inf_{T}\mathcal{L}(f,T),$ (3)
@@
$\mathcal{L}(f,T)=\int_{\mathcal{Y}}f(y)\,d\nu(y)+\int_{\mathcal{X}}\left(k\left(x,T(x,\cdot)\#\omega\right)-\int_{\mathcal{Z}}f(T(x,z))\,d\omega(z)\right)d\mu(x).$ (4)
This formulation allows our neural approach to amortize computation across conditioning values, avoiding the prohibitive cost of solving separate OT problems for each condition. We note here that the results in Korotin et al. (2023) can be extended to weak costs, but they are out of the scope of this work.
The conditioned problem. We start from equation 4 to integrate the conditioning mechanism into the problem formulation, enabling applications like global sensitivity analysis to compute transport-based indices efficiently across multiple conditioning values with a single learned model. We assume that the condition $c$ belongs to a measure space $\mathcal{C}$. From the theoretical perspective, the extension is straightforward:
@@
$\mathcal{L}(f,T,c)=\int_{\mathcal{Y}}f(y,c)\,d\nu(y)+\int_{\mathcal{X}}\left(k\left(x,T(x,\cdot,c)\#\omega\right)-\int_{\mathcal{Z}}f(T(x,z,c),c)\,d\omega(z)\right)d\mu(x).$ (6)